# This work was supported in part by Hong Kong Innovation and Technology Commission Funding Administrative System II (ITF Ref. No. GHP/110/20GD) and the National Natural Science Foundation of China (62192714, U21B2006).(Corresponding author: Hongwei Liu; Chang Gao.)

### v2

@@ Abstract
Multi-frame detection algorithms can effectively utilize the correlation between consecutive echoes to improve the detection performance of weak targets. Existing efficient multi-frame detection algorithms are typically based on three sequential steps: plot extraction via a relative low primary threshold, track search and track detection. However, these three-stage processing algorithms may result in a notable loss of detection performance and do not fully leverage the available echo information across frames. As to applying graph neural networks in multi-frame detection, the algorithms are primarily based on node classification tasks, which cannot directly output target tracks. In this paper, we reformulate the multi-frame detection problem as a link prediction task in graphs. First, we perform a rough association of multi-frame observations that exceed the low threshold to construct observation association graphs. Subsequently, the Multi-feature Link Prediction Network is designed based on graph neural networks, which integrates multi-dimensional information, including echo structure, Doppler information, and spatio-temporal coupling of plots. By leveraging the principle of link prediction, we unifies the processes of track search and track detection into one step to reduce performance loss and directly output target tracks. Experimental results indicate that, compared with traditional single-frame and multi-frame detection algorithms, the proposed algorithm improves the detection performance of weak targets while suppressing false alarms. Additionally, interpretable analysis shows that the designed network effectively integrates the utilized features, allowing for accurate associations between targets and false alarms.
@@
Multi-frame Detection, Graph Neural Network, Link Prediction, Weak Target Detection
@@ 1 INTRODUCTION
Target detection aims to determine whether a signal received by sensors originates from a target based on the signal strength[1]. Effective target detection serves as the foundation for subsequent operations such as target tracking and target recognition. The reliable detection of weak targets still remains a critical challenge that urgently needs to be addressed in systems such as radar and sonar. The traditional signal processing framework, which follows a sequential pattern of track after detection, struggles to balance detection performance with system processing load, and its drawbacks are particularly evident in weak target detection scenarios[2],[3],[4]. To address the shortcomings of traditional signal processing, multi-frame detection (MFD) methods leverage the regularity of target motion and the accumulation of target energy to correctly integrate the echo information along the unknown target trajectories in multiple frames, thereby suppressing false alarms while improving the detection performance for low signal-to-noise ratio (SNR) targets[5],[6],[7].
@@
Traditional MFD methods enable the successful detection of weak targets by directly utilizing un-thresholded raw echo data, but they exhibit significant drawbacks in high-resolution detection scenarios that require a large number of detection units[8],[9],[10],[11],[12]. On one hand, the large volume of raw observations will impose a substantial computational burden on the system. On the other hand, the observation uncertainty introduced by excessive false alarms makes it difficult to correctly associate targets. Although some methods focus on improving the accuracy of target association in different scenarios, they still do not consider the possibility of pre-filtering the candidate detection set based on the strength of the received echoes[13],[14],[15],[16],[17]. A viable approach to address the aforementioned challenge is a type of search-before-detect MFD method, which typically involves three steps[18]: First, a relatively low primary threshold is set for detection and plot extraction on each frame, which aims to maintain good detection performance for weak targets while filtering out observations unlikely to originate from actual targets, thereby reducing computational load. Second, for multi-frame observations that exceed the low threshold, potential tracks are searched by leveraging target kinematic constraints. Finally, a fusion decision is made along the potential tracks based on information such as echo energy. This kind of method transforms single-frame plot-level detection into multi-frame track-level detection, ensuring robust detection performance for weak targets without requiring successful detection in every frame. However, such sequential track search and track detection limit the overall performance by the effectiveness of the potential track search.
@@
Traditional track search methods primarily rely on position and Doppler information to establish kinematic constraints[18],[19],[20],[21]. In MFD methods that incorporate neural networks, reference [22] adopts a search-before-detect approach, first performing track search based on the coupled position and Doppler measurements, then conducting track detection using a data-driven method that integrates multi-dimensional information[23]. Reference [24] extends the scenario to bistatic radar systems based on this concept. Nevertheless, the model-driven track search used in these methods often lead to target missed associations when there is a model mismatch. Because it is difficult to apply unified constraint parameters that can reliably associate targets with different kinematic features, which in turn hinders the track detection process.
@@
Graphs, as a mathematical language in the physical world, can reasonably describe multi-frame data[25]. Compared to other neural network, graph neural networks (GNNs) are capable of modeling the data characteristics from both structural and functional perspectives, successfully extracting data features even with a limited number of labels[26],[27],[28],[29]. Although the sparse distribution of targets in the observation space results in a limited number of target labels, the use of GNNs can effectively address this shortcoming. In the detection methods which apply GNNs, reference [30] introduces a novel approach using GNNs to develop a message-passing solution for the inference task of massive multiple-input multiple-output (MIMO) detection in wireless communication. Reference [31] constructs graphs from point cloud data obtained from sensors in the field of autonomous driving, leveraging GNNs to fully exploit point features and relationship characteristics among points, achieving object detection and classification through a node prediction task. Reference [32], in a maritime scenario, similarly employs the node classification concept within GNNs to distinguish targets and clutter in multi-frame observations. However, this method cannot directly output target tracks based on node classification task and fails to incorporate domain knowledge to construct the features of edges.
@@
To overcome the limitations of model-driven track search and achieve target detection in environments with dense false alarms, a data-driven approach should enhance the accuracy of target association [33]. The process of track detection from multi-frame observations can be viewed as a nonlinear mapping from multi-dimensional information to the determination of whether a trajectory originates from a target. Furthermore, the track search and detection process is fundamentally equivalent to a clustering problem of spatio-temporal plot associations. Unlike the node classification in GNNs, link prediction within graphs can be utilized to integrate multi-dimensional features on different attributes, thereby accomplishing track search and detection in a unified manner[34],[35]. This approach allows for direct output of target trajectories from multi-frame observations, simplifying the processing logic while reducing the loss in track search, ultimately leading to improved target detection performance.
@@
Based on the above considerations, we proposes a multi-frame detection algorithm via graph neural network-based link prediction (GLP-MFD). The proposed algorithm first leverages prior knowledge of target kinematic to perform a rough association of consecutive multi-frame observations that exceed the low threshold, aiming to minimize target missed associations while constructing observation association graphs. The features like echo structure and Doppler information from observations are represented in nodes, while the spatio-temporal coupling relationship among multi-frame observations is represented in edges. Then the Multi-feature Link Prediction Network (MFLPN) is designed to carry out the multi-frame detection process within observation association graphs. Our contributions are summarized as follows.
@@
    By modeling the multi-frame detection as a link prediction task in graphs, we overcome the performance bottleneck in traditional MFD methods, which arises from model mismatches and insufficient utilization of echo information during the sequential track search and detection. By constructing multi-frame observations as observation association graphs, we perform a link prediction task to directly output target tracks, which enables integrated track search and detection, ensuring effective suppression of false alarms while achieving superior target detection performance.
@@
    By integrating information such as echo structures, Doppler measurements, and spatio-temporal coupling, we specifically design the Multi-feature Link Prediction Network which effectively exploits multi-dimensional features and leverage a unified network cost function to enable the network to perform link prediction tasks with greater accuracy.
@@
    We construct edge features based on coupled position and Doppler measurements. A novel message passing function is designed to integrate edge features while incorporating an attention mechanism. This approach effectively extracts spatio-temporal coupling information among plots, thereby enhancing the network performance.
The rest of this article is organized as follows. Section II briefly describes the MFD problem. Sections III and IV introduce the design of the GLP-MFD and its corresponding computational complexity analysis, respectively. Section V presents numerical results to validate the performance improvements of the proposed algorithm. Finally, Section VI concludes this article.
@@ 2 PROBLEM FORMULATION
Coherent radar systems can improve the target SNR through pulse compression and coherent integration, while obtaining Doppler measurements. After detection with the primary threshold and plot extraction, the $k$-th observation at frame $n$ can be represented as a vector:
@@
where $t_{k,n}$ is the time instant when the observation is taken, $r_{k,n}$ is the range measurement, $\theta_{k,n}$ is the azimuth measurement, $v_{k,n}$ is the potentially ambiguous radial velocity, $\boldsymbol{a}_{k,n}$ is the range-Doppler map and $\top$ denotes the transpose operation. For better readability, we considers a 2-D radar scenario, but the proposed algorithm can be easily extended to 3-D radar scenarios that include elevation information. The relationship between the true target state in Cartesian coordinates $\left(x_{k,n},\dot{x}_{k,n},y_{k,n},\dot{y}_{k,n}\right)$ and the radar measurements is as follows:
$r_{k,n}$ $=\sqrt{x_{k,n}^{2}+y_{k,n}^{2}}+w_{k,n}^{r}$ (2)
$\theta_{k,n}$ $=\arctan\frac{y_{k,n}}{x_{k,n}}+w_{k,n}^{\theta}$
@@
where $v_{u}$ denotes the range of the unambiguous radial velocity. $w_{k,n}^{r}$, $w_{k,n}^{\theta}$ and $w_{k,n}^{v}$ represent the noise of range, azimuth and radial velocity measurements, respectively. All three are zero-mean additive Gaussian white noise, with standard deviations proportional to the range, azimuth and velocity resolution, and inversely proportional to the square root of the SNR. Considering range ambiguity can be resolved by using techniques such as frequency stepping or transmitting orthogonal waveforms, we only focus on the potential ambiguity in radial velocity. The observation list corresponds to $D_{n}$ plots obtained in frame $n$ is defined as follows:
@@
The task of multi-frame target detection is to determine whether a target exists within the multi-frame observations $\left\{\mathbf{Z}_{l}\right\}_{l=n}^{n+L-1}$, where $L$ represents the number of frames processed jointly, also known as the length of processing sliding window. The operating scheme of the proposed GLP-MFD algorithm is depicted in Fig. 1. First, multi-frame observations that exceed the primary threshold $\gamma_{1}$ are roughly associated to construct observation association graphs. Subsequently, MFLPN is employed to complete the integrated track search and track detection process within the observation association graphs, after which the confidence score for each candidate track is obtained. The candidate track with a confidence score exceeding $\gamma_{2}$ will be declared as the potential track. Finally, confirmed target tracks are output after track pruning, which can be used for subsequent operations such as target tracking. The primary threshold $\gamma_{1}$ determines the optimal detection performance of the proposed scheme.
@@
To enhance the detection capability for weak targets, $\gamma_{1}$ is typically set to a relatively low value. Additionally, it is essential to minimize false alarms to prevent imposing an undue computational burden on the system.
@@ 3 MULTI-FRAME DETECTION VIA GRAPH NEURAL NETWORK-BASED LINK PREDICTION
The proposed GLP-MFD algorithm first utilizes prior knowledge of target kinematics and employs relaxed constraints tailored to the target to perform a rough association of multi-frame observations that exceed the primary threshold, thereby forming observation association graphs. This approach enables target plots to readily meet the constraints, ensuring successful associations and preventing performance degradation due to association failures. Then, we design MFLPN that integrates track search and track detection. Specifically, the Node Feature Embedding Network (NFEN) is developed to extract plot features such as echo structure and Doppler information by leveraging the strong fitting capabilities of Deep Neural Networks (DNNs) for nonlinear mapping[36],[37]. Subsequently, we design the Graph Attention Network with Spatio-Temporal Coupling Encoded in Edge Features (STEF-GAN) based on GNNs. By incorporating an attention mechanism and constructing edge features during message passing, this network more adequately extracts relational information that embody physical motion patterns among target plots. Finally, we develop an Observation Association Judgment Network (OAJN) to utilize the link prediction concept in order to output potential tracks, which will be declared as confirm tracks after pruning.
By employing the aforementioned data-driven approach to learn the complex nonlinear mapping from multi-frame information to target detection, the method integrates the two sequential steps into one, thereby simplifying the processing logic while reducing loss. Furthermore, it effectively leverages multi-dimensional features on different attributes, such as echo structure and spatio-temporal coupling information, resulting in superior performance in detecting weak targets while effectively suppressing false tracks.
@@ 3.1 Construction of observation association graphs
The target plots exhibit a certain spatio-temporal correlation across frames, allowing for using prior knowledge of target kinematics to roughly associate multi-frame observations that exceed a relative low primary threshold. An undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, referred to as the observation association graph is constructed, where $\mathcal{V}$ denotes the set of nodes representing multi-frame observations and $\mathcal{E}$ signifies the set of edges that connect successfully associated nodes. Correct target tracks exist within the graph, and we aim to accurately finding target tracks while eliminating false tracks. Due to the limited computational power of the system, it is unreasonable to associate all observations across frames and instead, multi-frame observations should be associated based on a certain constraint. To mitigate performance loss due to target missed associations, the selected constraint should be adaptable to targets. Since the velocities of the targets of interest typically have an upper bound, we leverage the velocity limit to ensure successful target association while roughly excluding false observations unlikely to be associated with targets, thereby reducing computational burden. Therefore, observation association graphs are constructed by using the maximum velocity constraint.
For two observations $\left(t_{1},r_{1},\theta_{1},v_{1}\right)$ and $\left(t_{2},r_{2},\theta_{2},v_{2}\right)$ from different frames, where $t_{1}<t_{2}$, the maximum velocity constraint is as follows:
$\|\mathbf{p}_{2}$ $-\mathbf{p}_{1}\|\leq v_{\max}(t_{2}-t_{1})$ (4)
$\mathbf{p}_{1}$ $=\left(r_{1}\cos\theta_{1},r_{1}\sin\theta_{1}\right)$
$\mathbf{p}_{2}$ $=\left(r_{2}\cos\theta_{2},r_{2}\sin\theta_{2}\right)\text{.}$
@@
The observations in multi-frame data $\left\{\boldsymbol{Z}_{l}\right\}_{l=n}^{n+L-1}$ that satisfy the maximum velocity constraint will be associated. Due to the difficulty in reliably detecting low SNR targets in every frame, and to ensure successful target association and track continuity, we stipulate that observations need to undergo association testing with those in subsequent $Q$ frames. Correspondingly, the maximum number of consecutive missed detections allowed for a track is set to $Q-1$. If a track has $Q$ consecutive frames without observations, it is deemed to be of poor quality and should be discarded. A sample of the constructed observation association graph is illustrated in Fig. 2, demonstrating the effective associations among target plots, which helps to avoid missed associations due to model mismatches. Additionally, the number of plots in the confirmed track is set to be at least $M$. In other words, any path in the graph with $M$ of more nodes is considered as a candidate track. And the link prediction acts on these candidate tracks to eliminate associations containing false alarms while preserving the associations between targets, thereby achieving integrated track search and track detection.

### v3

# {textblock*}

@@ Abstract
Multi-frame detection algorithms can effectively utilize the correlation between consecutive echoes to improve the detection performance of weak targets. Existing efficient multi-frame detection algorithms are typically based on three sequential steps: plot extraction via a relatively low primary threshold, track search, and track detection. However, these three-stage processing algorithms may result in a notable loss of detection performance and do not fully leverage the available echo information across frames. For the application of graph neural networks in multi-frame detection, the algorithms are primarily based on node classification tasks, which do not focus on directly outputting target tracks. In this paper, we reformulate the multi-frame detection problem as a link prediction task in graphs. First, we perform a rough association of multi-frame observations that exceed the primary threshold to construct observation association graphs. Subsequently, the multi-feature link prediction network is designed based on graph neural networks, which integrates multi-dimensional information, such as Doppler, signal structures, and spatio-temporal coupling of observations. By leveraging the principle of link prediction, we unify the process of track search and track detection into one step to reduce performance loss and directly output target tracks. Numerical results indicate that the proposed algorithm improves the detection performance of weak targets while suppressing false alarms, compared with traditional single-frame and multi-frame detection algorithms. Additionally, interpretability analysis shows that the designed network effectively integrates the utilized features, allowing for accurate target associations.
@@
Multi-frame Detection, Graph Neural Network, Link Prediction, Weak Target Detection.
@@ 1 INTRODUCTION
Target detection aims to determine whether a signal received by sensors originates from a target based on the signal strength [1]. Effective target detection serves as the foundation for subsequent operations such as target tracking and target recognition. The reliable detection of weak targets still remains a critical challenge that urgently needs to be addressed in systems such as radar and sonar. The traditional signal processing framework, which follows a sequential pattern of track after detection, struggles to balance detection performance with system processing load [2],[3],[4]. Its drawbacks are particularly evident in weak target detection scenarios. To address the shortcomings of traditional signal processing, multi-frame detection (MFD) methods leverage the regularity of target motion and the accumulation of target energy to improve the detection performance for low signal-to-noise ratio (SNR) targets [5],[6],[7],[8].
@@
Traditional MFD methods enable the successful detection of weak targets by directly utilizing un-thresholded raw data [9],[10],[11],[12],[13]. But they exhibit drawbacks in high-resolution detection scenarios that require a large number of detection cells. On one hand, the large volume of observations will impose a substantial computational burden on the system. On the other hand, the observation uncertainty introduced by excessive false alarms makes it difficult to perform accurate target associations. Although some methods focus on improving the accuracy of target measurements in different scenarios, they still do not consider pre-filtering the candidate detection set based on the strength of the received echoes [14],[15],[16]. A viable approach to address the aforementioned challenge is a type of three-stage MFD methods, which typically involves three steps [17],[18],[19]: First, a relatively low primary threshold is set for detection and plot extraction on each frame. The low threshold aims to maintain good detection performance for weak targets while preliminarily filtering out observations unlikely to originate from targets, thereby reducing computational load. Second, for multi-frame observations that exceed the primary threshold, potential tracks are searched by leveraging target kinematic constraints. Finally, a fusion detection is made along the potential tracks based on information such as echo energy. This kind of methods transforms single-frame plot-level detection into multi-frame track-level detection, improving detection performance for weak targets. However, such sequential track search and track detection hinder the overall performance by the limitations of the potential track search.
@@
Traditional methods of track search primarily rely on position and Doppler information to establish kinematic constraints [18],[20],[21],[22]. In MFD methods that incorporate neural networks, reference [23] adopts the three-stage scheme, first performing model-driven track search based on the position and Doppler measurements, then conducting track detection using a data-driven method [24]. Reference [25] extends the scenario to the bistatic radar system based on this concept. Nevertheless, the model-driven track search used in these methods often leads to target missed associations when there is a model mismatch. Because it is difficult to apply unified constraint parameters to reliably associate targets with different kinematic features, which in turn hinders the track detection process.
@@
Graphs, as a mathematical language in the physical world, can reasonably describe multi-frame data [26]. Compared to other neural networks, graph neural networks (GNNs) are capable of modeling data characteristics from both structural and functional perspectives, successfully extracting data features even with a limited number of labels [27],[28],[29]. Although the sparse distribution of targets in the observation space results in a limited number of target labels, the use of GNNs can effectively address this issue. In the detection methods which apply GNNs, reference [30] introduces a novel approach using GNNs to develop a message-passing solution for the inference task of massive multiple-input multiple-output (MIMO) detection in wireless communication. Reference [31] constructs graphs from point cloud data obtained from sensors in the field of autonomous driving. The method leverages GNNs to fully exploit point features and relationship characteristics among points, achieving object detection and semantic segmentation through a node classification task. Reference [32], in a maritime scenario, similarly employs the node classification concept within GNNs to distinguish targets and clutter in multi-frame observations. However, the utilization of the node classification task does not focus on directly outputting target tracks. Additionally, edge features can be considered for design by incorporating domain knowledge.
@@
To overcome the limitations of model-driven track search and achieve target detection in environments with dense false alarms, a data-driven approach should enhance the accuracy of target associations [33]. The process of track detection from multi-frame observations can be viewed as a nonlinear mapping from multi-dimensional information to the determination of whether a trajectory originates from a target. Furthermore, the track search and detection process is fundamentally equivalent to a clustering problem of spatio-temporal plot associations. Unlike the node classification, link prediction within graphs can be utilized to integrate multi-dimensional information on different attributes to accomplish track search and track detection in a unified manner [34],[35]. This approach allows for direct output of target trajectories from multi-frame observations, simplifying the processing logic while reducing the loss in track search, ultimately leading to improved target detection performance.
@@
Based on the above considerations, we propose a multi-frame detection algorithm via graph neural network-based link prediction (GLP-MFD). The proposed algorithm first performs a rough association for multi-frame observations that exceed the primary threshold. And the successfully associated observations are constructed as observation association graphs. The features like Doppler information and signal structures of observations are represented in nodes, while the spatio-temporal coupling relationships among multi-frame observations are represented in edges. Then the multi-feature link prediction network (MFLPN) is designed to carry out the multi-frame detection process within observation association graphs. Our contributions are summarized as follows.
@@
    By modeling the multi-frame detection as a link prediction task in graphs, we overcome the performance bottleneck in traditional MFD methods, which arises from model mismatches and insufficient utilization of echo information during sequential track search and track detection. By constructing multi-frame observations as observation association graphs, we perform a link prediction task to directly output target tracks. The proposed algorithm enables integrated track search and track detection, achieving superior target detection performance while ensuring effective suppression of false alarms.
@@
    By integrating information such as Doppler, signal structures, and spatio-temporal coupling, we specifically design MFLPN which effectively exploits multi-dimensional features. Additionally, we leverage a unified cost function to enable the network to perform the link prediction task with higher accuracy.
@@
    We design edge features based on coupled position and Doppler measurements. A novel message passing function is designed to integrate edge features while incorporating an attention mechanism. This approach effectively extracts spatio-temporal coupling information among observations, thereby enhancing the network performance.
The rest of this article is organized as follows. Section 2 briefly describes the MFD problem. Sections 3 and 4 introduce the design of GLP-MFD and its corresponding computational complexity analysis, respectively. Section 5 presents numerical results to validate the performance improvement of the proposed algorithm. Finally, Section 6 concludes this article.
@@ 2 PROBLEM FORMULATION
Coherent systems can improve the target SNR through pulse compression and coherent integration, while obtaining Doppler measurements. After detection and plot extraction with the primary threshold $\gamma_{1}$, the $k$-th observation at frame $n$ can be represented as a vector:
@@
where $t_{k,n}$ is the time instant when the observation is taken, $r_{k,n}$ is the range measurement, $\theta_{k,n}$ is the azimuth measurement, $v_{k,n}$ is the potentially ambiguous radial velocity, $\boldsymbol{a}_{k,n}$ is the range-Doppler map and $\top$ denotes the transpose operation. For better readability, we consider a 2-D scenario. The relationship between the true target state $\left(x_{k,n},\dot{x}_{k,n},y_{k,n},\dot{y}_{k,n}\right)$ in the Cartesian coordinate and the sensor measurements is as follows:
$r_{k,n}$ $=\sqrt{x_{k,n}^{2}+y_{k,n}^{2}}+w_{k,n}^{r}\text{,}$ (2)
$\theta_{k,n}$ $=\arctan\frac{y_{k,n}}{x_{k,n}}+w_{k,n}^{\theta}\text{,}$
@@
where $v_{u}$ denotes the range of the unambiguous radial velocity. $w_{k,n}^{r}$, $w_{k,n}^{\theta}$ and $w_{k,n}^{v}$ represent the noise of range, azimuth and radial velocity measurements, respectively. All three are zero-mean additive Gaussian white noise, with standard deviations proportional to the range, azimuth and velocity resolution, and inversely proportional to the square root of SNR. Considering range ambiguity can be resolved by using techniques such as frequency stepping or transmitting orthogonal waveforms, we only focus on the potential ambiguity in radial velocity. The observation list corresponding to $D_{n}$ plots obtained in frame $n$ is defined as follows:
@@
The task of multi-frame target detection is to determine whether a target exists within the multi-frame observations $\left\{\mathbf{Z}_{l}\right\}_{l=n}^{n+L-1}$, where $L$ represents the number of frames processed jointly, also known as the length of the processing sliding window. In traditional three-stage MFD methods, a model-driven track search is first preformed based on target kinematic constraints:
$\Lambda\left(\mathbf{Z}_{n},\mathbf{Z}_{n+1},...,\mathbf{Z}_{n+L-1}\right)\leq\boldsymbol{\mathit{\Upsilon}}\text{,}$ (4)
where $\Lambda$ represents the function of constraints and $\boldsymbol{\mathit{\Upsilon}}$ represents the vector of constraint parameters. The multi-frame observations meeting the constraints will be associated to form potential target tracks. And those failing to meet the constraints will be regarded as false alarms. Then, a detector for potential tracks can be designed using methods such as the generalized likelihood ratio test [36],[37],[38],[39]. The potential tracks with test statistic exceeding the final detection threshold $\gamma_{2}$ will be declared as confirm tracks after pruning. However, such model-driven track search usually suffers from a performance loss due to target missed associations. And the detector is hard to derive when fusing multi-dimensional information on different attributes. Therefore, a data-driven method is proposed in this work to address these problems.
@@ 3 MULTI-FRAME DETECTION VIA GRAPH NEURAL NETWORK-BASED LINK PREDICTION
The operating scheme of the proposed GLP-MFD algorithm is depicted in Fig. 1. First, detection and plot extraction are performed on the raw data using the primary threshold $\gamma_{1}$. To enhance the detection capability for weak targets, $\gamma_{1}$ is typically set to a relatively low value. Additionally, it is essential to preliminarily eliminate false alarms to prevent imposing an undue computational burden on the system. Subsequently, multi-frame observations are roughly associated by applying loose constraints to construct observation association graphs. This step aims to enable target observations to readily meet the constraints, ensuring successful associations and preventing performance degradation due to association failure. Then, the designed MFLPN is employed to complete the integrated track search and track detection process within the observation association graphs. The potential tracks can be obtained after the network processing with the final detection threshold $\gamma_{2}$. Finally, confirmed tracks are output after track pruning, which can be used for subsequent operations such as target tracking.
By employing the aforementioned data-driven approach, the designed network effectively learns the complex nonlinear mapping from multi-frame information to target detection. In addition, the proposed algorithm integrates the two sequential steps of track search and track detection into one, thereby simplifying the processing logic while reducing loss. Furthermore, it effectively leverages multi-dimensional information on different attributes, leading to superior performance in detecting weak targets.
@@ 3.1 Construction of observation association graphs
An undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, referred to as the observation association graph is constructed by using a loose constraint. $\mathcal{V}$ denotes the set of nodes representing multi-frame observations and $\mathcal{E}$ signifies the set of edges that connect successfully associated nodes. Correct target tracks exist within the graph and we aim to accurately find target tracks while eliminating false tracks. Due to the limited computational power of the system, it is unreasonable to associate all observations across frames. The velocities of the targets of interest typically have an upper bound. We leverage the velocity limit to ensure successful target associations while roughly excluding false observations unlikely to be associated with targets. In this way, we mitigate performance loss due to target missed associations and reduce the computational burden. Therefore, observation association graphs are constructed by using the maximum velocity constraint. For two observations $\left(t_{1},r_{1},\theta_{1},v_{1}\right)$ and $\left(t_{2},r_{2},\theta_{2},v_{2}\right)$ from different frames, where $t_{1}<t_{2}$, the maximum velocity constraint is as follows:
$\|\mathbf{p}_{2}$ $-\mathbf{p}_{1}\|\leq v_{\max}(t_{2}-t_{1})\text{,}$ (5)
$\mathbf{p}_{1}$ $=\left(r_{1}\cos\theta_{1},r_{1}\sin\theta_{1}\right)\text{,}$
$\mathbf{p}_{2}$ $=\left(r_{2}\cos\theta_{2},r_{2}\sin\theta_{2}\right)\text{,}$
@@
Considering the difficulty in reliably detecting weak targets in each frame, as well as the need to ensure successful target associations and track continuity, we stipulate that observations need to undergo association testing with those in subsequent $Q$ frames. Correspondingly, the maximum number of consecutive missed observations allowed for a track is set to $Q-1$. If a track has $Q$ or more consecutive frames without observations, it is deemed to be of poor quality and should be discarded. An illustration of the constructed observation association graph is shown in Fig. 2. The nodes represent multi-frame observations which exceed the primary threshold. The edges represent observations undergoing association testing that satisfy the maximum velocity constraint. The illustration demonstrates the successful associations among target observations, which helps to avoid performance loss due to missed associations. Additionally, the number of observations in the confirmed track is set to be at least $M$. In other words, any path in the graph with $M$ or more nodes is considered as a candidate track. And the link prediction acts on these candidate tracks to eliminate associations containing false alarms while preserving the associations between targets, thereby achieving integrated track search and track detection.